# CEFR-Annotated WordNet: LLM-Based Proficiency-Guided Semantic Database for Language Learning

### v1

@@ 10. Limitations
A key limitation of our work is the need to expand the coverage of CEFR-level annotations in WordNet. Despite our novel automated pipeline, we annotated only 10,644 senses, which represents approximately 5% of WordNet’s total. This low coverage is a direct result of the limited availability of sense-level CEFR labels in the EVP Online, a foundational resource for many dictionaries that constrains further expansion. To mitigate this, we developed lexical CEFR-level classifiers trained on our large, sense-annotated corpus, which achieved a maximum Macro-F1 score of 0.81. While these classifiers can predict levels for previously unseen senses from their usage context, they are currently less accurate than our primary gloss-based method. Consequently, improving their accuracy, capability, and robustness is essential for reliable, large-scale applications. In parallel, it is crucial to quantify the impact of annotation errors on L2 learners and establish acceptable error thresholds for educational use. Another limitation is that our evaluation targeted only single words, whereas related work in tasks such as LCP and models like ME6 Contextual also includes MWEs. Our preliminary analysis indicates that most MWEs map to a single CEFR level, suggesting that our knowledge-based approach could classify them with high accuracy. However, significant challenges remain, particularly in identifying implicit MWEs within texts (such as in CALL systems analyzing textbooks) or handling MWEs that are not present in the EVP. Addressing these limitations will require the development of more powerful and context-aware classifiers.
@@ 11. Bibliographical References
- Vicente et al. (2014) Iñaki San Vicente, Rodrigo Agerri, and German Rigau. 2014. Simple, robust and (almost) unsupervised generation of polarity lexicons for multiple languages. In Proceedings of the 14th Conference of the European Chapter of the Association for Computational Linguistics (EACL 2014), pages 88–97.
@@ 13.1. Parameters and Prompts
To measure the semantic similarity between and , we used GPT-4o (checkpoint gpt-4o-2024-08-06) with the prompt shown in Figure 3. The system content was set to “You are a professional linguist,” and the temperature was set to 0 to ensure deterministic outputs.
@@
We used GPT‑5 (checkpoint gpt-5-2025-08-07) as the base model for our classifiers, setting the system content to “You are an expert rater for the Common European Framework of Reference for Languages (CEFR).” and the parameter reasoning_effort to “high.” Figures 4 and 5 show the prompt templates used for the zero-shot and few-shot LLM classifiers, respectively. In a preliminary experiment, we provided the LLMs with full descriptions of the CEFR levels based on the official can-do descriptors. However, we observed no significant difference in classification performance compared to using the simplified descriptions shown in the figures. Therefore, for efficiency, we opted for the prompts with simplified descriptions in our experiments.
@@ 13.2. Confusion Matrices
where denotes the number of target words with the actual CEFR level and is the number of those words classified as level , i.e., . The diagonal elements correspond to the recall for each level; thus, higher values along the diagonal indicate greater accuracy. Because the CEFR levels are ordinal, misclassifications that fall closer to the diagonal (i.e., to an adjacent level) are less disruptive for language learners.
The ME6 Contextual models achieve high recall for the lower levels (A1 and A2) and the highest level (C2). However, as shown in Figures 6 and 6, when the model is trained on either the EVP or SemCor-CEFR corpus alone, errors at the intermediate and advanced levels (B1–C2) are more broadly distributed. By contrast, combining both resources (Figure 6) narrows the distribution of these errors, with most misclassifications occurring between adjacent levels. This result underscores the benefit of leveraging both resources jointly.
As shown in Figure 6, the zero-shot LLM achieves high recall for A1 (82.8%) and moderate recall for A2 (58.5%), but its performance declines for B1 (36.9%) and is notably poor for C1 and C2 (26.0% and 12.1%, respectively). The model tends to misclassify advanced-level senses as B2 (e.g., 54.2% of C1 and 44.2% of C2), indicating a tendency to collapse more challenging senses into an intermediate level. Few-shot prompting (Figures 6 and 6) partially mitigates this issue, improving recall for C1 and C2. However, recall for A2 declines compared to that with the zero-shot baseline, and B1 performance remains comparable, indicating that the gains from few-shot prompting are uneven across proficiency levels.
By contrast, the FT and FT+KB models substantially improve performance across all CEFR levels. When fine-tuned on the Mixture dataset (Figures 6 and 6), the FT model attains recall above 70% for levels A1–B2 and just under 70% for C2. The FT+KB model improves recall further, surpassing 80% for B1–B2, reaching approximately 91% for A1, remaining in the high-70% range for A2 and C2, and around 70% for C1. Moreover, the confusion matrices for these models show that errors are concentrated near the diagonal, indicating that they typically involve adjacent CEFR levels, which minimizes pedagogical disruption. Despite these gains, C1 remains a challenging level, and C2 instances are often misclassified as B2. This pattern persists even when fine-tuning on the individual EVP or SemCor-CEFR corpora (Figures 6 and 6), suggesting the issue is not an artifact of our annotation method. Instead, it likely stems from the CEFR-level distribution in the EVP data and the properties of the fine-tuning process. Addressing these residual errors will require further investigation.

### v2

@@ 10. Limitations
A key limitation of this work is the limited coverage of CEFR-level annotations in WordNet. Despite the automated pipeline, only 10,644 senses were annotated, representing approximately 5% of the full inventory. This limited coverage stems from the restricted availability of sense-level CEFR labels in the EVP Online, which serves as the primary source for alignment and constrains further expansion. Beyond the EVP Online, several CEFR-related lexical resources are available, such as CEFRLex and SeCoDa, as well as lexical complexity datasets including CompLex 2.0. However, these resources differ in granularity and label space. Proficiency information is often provided at the word level rather than the sense level, sense inventories are not aligned with WordNet, and labels are not always expressed as CEFR levels. As a result, they cannot be directly integrated into our sense-level annotations without additional mapping. Developing principled methods to use them as auxiliary supervision remains future work. To mitigate the limited coverage, we developed lexical CEFR-level classifiers trained on the large, sense-annotated corpus, achieving a maximum Macro-F1 score of 0.81. Although these classifiers can predict levels for previously unseen senses based on contextual usage, they are currently less accurate than the primary gloss-based transfer method. Improving their accuracy, robustness, and generalization is therefore essential for reliable large-scale deployment. In parallel, it is important to assess how annotation errors affect L2 learners and establish acceptable error thresholds for educational applications. Another limitation is that the evaluation focused exclusively on single words, whereas related work on LCP, including models such as ME6 Contextual, also considers multiword expressions. Preliminary analysis suggests that most MWEs correspond to a single CEFR level, indicating that the knowledge-based component could classify them with high accuracy. However, challenges remain in identifying implicit MWEs in running text, such as in CALL systems analyzing textbooks, and in handling expressions not covered by the EVP. Addressing these issues will require more advanced and context-sensitive classification methods.
@@ 11. Bibliographical References
- San Vicente et al. (2014) Iñaki San Vicente, Rodrigo Agerri, and German Rigau. 2014. Simple, robust and (almost) unsupervised generation of polarity lexicons for multiple languages. In Proceedings of the 14th Conference of the European Chapter of the Association for Computational Linguistics (EACL 2014), pages 88–97.
@@ 13.1. Parameters and Prompts
To ensure reproducibility, we report the exact checkpoint of the OpenAI model used in this study. Different checkpoints were used for gloss similarity measurements (Section 4) and classifier experiments (Section 5), as these components were conducted at different times. For each task, we used the most recent checkpoint available at the time of execution.
@@
To measure the semantic similarity between and , we used GPT-4o (checkpoint gpt-4o-2024-08-06) with the prompt shown in Figure 3. The system message was set at “You are a professional linguist,” and the temperature was set at 0 to ensure deterministic outputs.
@@
We used GPT-5 (checkpoint gpt-5-2025-08-07) as the base model for our classifiers. The system message was set at “You are an expert rater for the Common European Framework of Reference for Languages (CEFR).” and the parameter reasoning_effort was set at “high.” Figures 4 and 5 present the prompt templates used for the zero-shot and few-shot LLM classifiers, respectively. In a preliminary experiment, we provided the LLMs with full CEFR level descriptions based on the official can-do descriptors. However, we observed no significant difference in classification performance compared to using the simplified descriptions shown in the figures. For efficiency, we therefore used the prompts with simplified descriptions in our experiments.
@@
| # Distinct CEFR Labels | # Sense Keys | Share (%) |
| 1 | 10,308 | 96.84 |
| 2 | 321 | 3.02 |
| 3 | 15 | 0.14 |
| Row Level | Column Level | |||||
| A1 | A2 | B1 | B2 | C1 | C2 | |
| A1 | — | 33 | 31 | 13 | 8 | 2 |
| A2 | 33 | — | 43 | 24 | 4 | 6 |
| B1 | 31 | 43 | — | 86 | 12 | 13 |
| B2 | 13 | 24 | 86 | — | 24 | 40 |
| C1 | 8 | 4 | 12 | 24 | — | 27 |
| C2 | 2 | 6 | 13 | 40 | 27 | — |
| Sense Key | Lemma | PoS | CEFR (3 Labels) | WordNet Gloss (Definition) |
| bad%3:00:00:: | bad | Adjective | A1/A2/C1 | having undesirable or negative qualities |
| block%2:35:02:: | block | Verb | B2/C1/C2 | obstruct |
| close%2:41:00:: | close | Verb | A2/B2/C2 | cease to operate or cause to cease operating |
| dance%1:04:00:: | dance | Noun | A1/A2/B1 | taking a series of rhythmical steps (and movements) in time to music |
| find%2:39:02:: | find | Verb | A1/A2/B1 | discover or determine the existence, presence, or fact of |
| give%2:32:02:: | give | Verb | A1/A2/B1 | bestow |
| give%2:36:00:: | give | Verb | A1/A2/B1 | give or supply |
| hard%3:00:06:: | hard | Adjective | A1/B1/C1 | not easy |
| miss%2:32:00:: | miss | Verb | A2/B1/B2 | fail to experience |
| safe%3:00:01:: | safe | Adjective | A1/A2/B1 | free from danger or the risk of harm |
| schedule%1:10:00:: | schedule | Noun | A2/B1/B2 | an ordered list of times at which things are planned to occur |
| shake%2:29:00:: | shake | Verb | B1/B2/C2 | move with or as if with a tremor |
| start%2:36:01:: | start | Verb | A1/B1/B2 | get off the ground |
| start%2:38:00:: | start | Verb | A1/B1/B2 | begin or set in motion |
| start%2:38:01:: | start | Verb | A1/B1/B2 | get going or set in motion |
13.2. Multi-Labeled Sense Keys
In WordNet, a sense key is the canonical string identifier of an individual sense, encoding the lemma, PoS, and a sense index. Because our CEFR annotation transfers levels from EVP senses to WordNet senses via gloss alignment, a single WordNet sense key may occasionally receive multiple CEFR labels. This occurs when a WordNet gloss is sufficiently similar, with a similarity score of , to more than one EVP gloss and those EVP glosses carry different CEFR levels. Such multi-label assignments primarily reflect differences in gloss granularity between the two resources.
How frequent are multi-labeled sense keys?
Table 6 summarizes the number of distinct CEFR labels assigned to each WordNet sense key. The vast majority of sense keys are unambiguous: 10,308 sense keys (96.84%) receive exactly one label. Multi-labeled cases are rare, with 321 sense keys (3.02%) assigned two labels and only 15 sense keys (0.14%) assigned three labels. Overall, multi-label assignments affect approximately 3.16% of all annotated sense keys. This indicates that the conservative alignment threshold produces predominantly single-level annotations while retaining coverage for borderline cases.
Which level combinations co-occur?
To characterize the nature of multi-labeling, Table 7 reports pairwise co-occurrence counts of CEFR levels within the multi-labeled sense keys. Two-label cases contribute one pair, and three-label cases contribute three pairs, yielding 366 total co-occurrence pairs in Table 7. A clear pattern emerges. The most frequent co-occurrence is B1–B2 (86), followed by A2–B1 (43), B2–C2 (40), and A1–A2 (33). Overall, adjacent-level pairs (A1–A2, A2–B1, B1–B2, B2–C1, C1–C2) dominate. This pattern suggests that multi-labeling typically occurs near proficiency boundaries rather than arising from arbitrary mismatches. Less frequent and more distant pairs, such as A1–C1 and A2–C2, may reflect particularly broad WordNet glosses or EVP senses whose pedagogical sequencing differs substantially across sub-senses.
What do three-label cases look like?
Table 8 lists the 15 sense keys assigned three distinct CEFR labels. These cases are dominated by highly frequent and semantically broad lemmas (e.g., bad, close, find, give, start), which are associated with short and general WordNet glosses. Such senses can plausibly align with EVP sub-senses introduced at different stages, for example early concrete uses versus later, more abstract or specialized uses. Importantly, the extremely small number of three-label cases suggests that wide label dispersion is exceptional; most multi-labeled sense keys involve only two nearby levels.
@@ 13.3. Confusion Matrices
where denotes the number of target words with the actual CEFR level and is the number of those words classified as level , i.e., . The diagonal elements represent the recall for each level; higher values along the diagonal therefore indicate greater accuracy. Because the CEFR levels are ordinal, misclassifications that fall near the diagonal, that is, those assigned to adjacent levels, are less disruptive for language learners.
The ME6 Contextual models achieve high recall at the lower levels, A1 and A2, as well as at the highest level, C2. However, as shown in Figures 6 and 6, when the model is trained on either the EVP or SemCor-CEFR corpus alone, errors at the intermediate and advanced levels, B1 to C2, are more widely distributed. By contrast, combining both resources (Figure 6) reduces this dispersion, with most misclassifications occurring between adjacent levels. This result highlights the advantage of jointly leveraging both resources.
As shown in Figure 6, the zero-shot LLM achieves high recall for A1 (82.8%) and moderate recall for A2 (58.5%). Performance declines at B1 (36.9%) and is particularly poor at C1 and C2 (26.0% and 12.1%, respectively). The model frequently misclassifies advanced-level senses as B2, for example 54.2% of C1 and 44.2% of C2 cases, indicating a tendency to collapse more difficult senses into an intermediate level. Few-shot prompting (Figures 6 and 6) partially alleviates this issue by improving recall at C1 and C2. However, recall at A2 declines relative to the zero-shot baseline, and performance at B1 remains similar. These results indicate that the gains from few-shot prompting are uneven across proficiency levels.
By contrast, the FT and FT+KB models substantially improve performance across all CEFR levels. When fine-tuned on the Mixture dataset (Figures 6 and 6), the FT model achieves recall above 70% for levels A1–B2 and just below 70% for C2. The FT+KB model further improves recall, exceeding 80% for B1–B2, reaching approximately 91% for A1, remaining in the high-70% range for A2 and C2, and around 70% for C1. The corresponding confusion matrices show that errors are concentrated near the diagonal, meaning that most misclassifications occur between adjacent CEFR levels. This pattern reduces potential pedagogical disruption. Despite these improvements, C1 remains challenging, and C2 instances are often misclassified as B2. This pattern also appears when fine-tuning on the EVP or SemCor-CEFR corpora individually (Figures 6 and 6), suggesting that it is not an artifact of the annotation method. Rather, it likely reflects the CEFR-level distribution in the EVP data and characteristics of the fine-tuning process. Further investigation is needed to address these residual errors.